### 08 Feb 2024 v3

@@ 1 Introduction
Simulating wave propagation is an important task for many applications. For instance, when searching natural gas, wave propagation computation form the forward part of a numerical method for solving the inverse problem of geophysical inversion. By accurately analyzing the reflections and transmissions generated by complex media discontinuities, it becomes possible to characterize underground formations. However, traditional numerical computations often demand a computationally expensive fine grid representation to guarantee stability.
@@
$\mathfrak{u}_{n+1}\coloneqq\mathfrak{u}(x,t+\Delta t)=\mathcal{F}_{\Delta t}\mathfrak{u}_{n}\approx\Psi_{\Delta t}^{\theta}\mathfrak{u}_{n}\coloneqq\mathcal{I}^{\theta}\mathcal{G}_{\Delta t}\mathcal{R}^{0}.$ (1)
@@ 2 Finite-Difference-Based Wave Propagators
As the two solvers operate on different Cartesian grids, with $\delta x<\Delta x$ and $\delta t<\Delta t$, we define the restriction operator $\mathcal{R}$, which transforms functions from a fine to a coarse grid, and the prolongation operator $\mathcal{I}$, which maps the inverse relation. The enhanced variants consist of (a) bilinear interpolations denoted as $\mathcal{R}^{0}$ and $\mathcal{I}^{0}$, while $\mathcal{I}^{0}\mathcal{R}^{0}\mathfrak{u}\neq\mathfrak{u}$, and (b) neural network components denoted as $\mathcal{I}^{\theta}\equiv\Lambda^{{\dagger}}\mathcal{I}_{\Delta t^{\star}}^{~{}\theta}\Lambda$, while the lower index indicates that the neural networks are trained on the step size $\Delta t^{\star}$. For improved neural network inference, we use the transition operator $\Lambda$ to transform physical wave fields $(u,u_{t})$ to energy component representations $(\nabla u,c^{-2}u_{t})$, with $\Lambda^{{\dagger}}$ as the pseudo-inverse (see also energy1 ). Figure 1 provides a schema visualizing the wave argument transitions.
Figure 1: Detailed schematic of E2E-JNet3 (adapted from nguyen2022numerical ). Each convolutional block (blue) encompasses a 3x3 convolutional layer ($\text{groups}=3$, $\text{padding}=1$), followed by a batch normalization and a ReLU activation function. To reduce the resolution, the 3rd and 6th convolutional block (red) employs a stride of two. Connectivity within the network is depicted by arrows, with the dashed arrow specifically indicating a single application of $G_{\Delta t^{\star}}$.
@@
A simple model with bilinear interpolation (E2E-V, $\mathcal{I}^{0}\Psi_{\Delta t^{\star}}^{\theta}\mathcal{R}^{0}$) is used as a baseline. Each variant changes the baseline by exactly one aspect. This allows us to isolate the effect of each modification on the performance of the architecture. The four investigated end-to-end models $\Psi_{\Delta t^{\star}}^{\theta,0}\coloneqq\mathcal{I}_{\Delta t^{\star}}^{~{}\theta}G_{\Delta t^{\star}}\mathcal{R}^{0}$ are:
@@
The second baseline is taken from nguyen2022numerical and denoted as the modular, not end-to-end 3-level JNet (NE2E-JNet3), $\mathcal{I}_{\Delta t^{\star}}^{~{}\theta}(G_{\Delta t^{\star}}\mathcal{R}^{0}$), while results of $\mathcal{G}_{\Delta t^{\star}}$ are used to separately train the E2E-JNet3 upsampling component.
@@ 3 Data Generation Approaches
    with $x\in[-1,1]^{2}$ and $\frac{1}{\sigma^{2}}\sim\mathcal{N}(250,10)$. $\tau\in[-0.5,0.5]^{2}$ for $i\in\{1,2\}$ represents the displacement of the Gaussian pulse’s location from the center.
@@
We sample $c\in\mathfrak{F}_{c}$ from randomly chosen subregions of two synthetic geological structures, Marmousi marmousi and BP bp , that are mapped onto the spatial grid $h\mathbb{Z}^{2}\cap[-1,1]^{2}$. For testing, four manually modified media are added to examine rapid variations in velocity:
@@ 3.1 Multi-Step Training
$\mathfrak{u}_{n+k}\coloneqq\mathfrak{u}(t_{n}+k\Delta t)=(F_{\Delta t})^{k}\mathfrak{u}_{n}\approx(\Psi_{\Delta t}^{\theta})^{k}\mathfrak{u}_{n}.$ (5)
@@ 3.2 Parareal Algorithm
Identical to nguyen2022numerical , our implemented scheme iteratively refines the solution using the difference between $\mathcal{F}_{\Delta t}$ and $\mathcal{G}_{\Delta t}$ for each subinterval $\Delta t$. In particular, missing high-frequency components occur due to the transition to a lower grid resolution, or a too simple numerical algorithm. Therefore, a more elaborate model $\Psi_{\Delta t}^{\theta}$ is required for convergence. Formally, we rearrange Eq. 1 for the time stepping of $\mathcal{F}_{\Delta t}$, and replace $\mathcal{F}_{\Delta t}(\mathcal{I}\mathcal{R})\mathfrak{u}_{n}$ by a computationally cheaper strategy $\Psi_{\Delta t}^{\theta}$ end-to-end:
@@
Parareal iterations alter a given initial sequence of wave fields $\mathfrak{u}_{n}^{0}$ to $\mathfrak{u}_{n}^{k}$ for $n\in\mathbb{N}$. This means that neural operators should be trained to map $\mathfrak{u}_{n+1}^{k}$ to $\mathcal{F}_{\Delta t}\mathfrak{u}^{k}_{n}$. Therefore, appropriate training patterns for this setup would naturally differ from those found in $\mathcal{D}$, and the dataset for use with Parareal should be sampled from a different distribution, denoted as $\mathcal{D}^{p}$.
@@ 4 Evaluation Setup
A. Comprehensive Training ($\mathcal{D}^{p}_{\text{train}}$): The models are trained according to the Parareal scheme in Eq. 7 with $K=4$ using a random sample that constitutes a quarter of the original dataset $\mathcal{D}$ for fair comparisons.
@@ 5 Discussion
An overview of the upsampling architecture performances can be found in Table 1. As expected, the larger networks (E2E-Tira and E2E-Trans) performed slightly better compared to the 3-level JNet architecture, but for the ResNet architecture (E2E-JNet5), more weights did not necessarily increase accuracy. Consequently, we theorize that the design of the ResNet may be insufficient for capturing high-fidelity wave patterns, while especially highly-connected layers with an optimized feature and gradient flow (E2E-Tira) are better suited. Given that E2E-JNet3 ($\mathcal{D}^{w,m}$) had only a slightly worse average energy MSE on the test set, we generally advise against using the expensive models in our setup.
@@
While models trained with $\mathcal{D}^{p}_{\text{train}}$ appear to have an unstable training progress, applying E2E-JNet3 ($\mathcal{D}^{p}_{\text{refine}}$) within the Parareal scheme showed better accuracy than E2E-JNet3 with Parareal (cf. Figure 5). As this training method improved the stability of Parareal, sampling the causality of concurrently solving multiple time intervals is an efficient enhancement to our end-to-end structure.
@@ 6 Conclusion
All end-to-end variants without training modifications outperformed the modular framework of nguyen2022numerical . In particular, the lightweight end-to-end 3-level JNet (E2E-JNet3) performed reasonably well given its low computation cost, and was further improved through a weighted, multi-step training scheme ($\mathcal{D}^{w,m}$) to feature time-dependent wave dynamics without adding complexity to the model or substantially extending the training duration. Similarly, the Parareal iterations using the neural propagator trained by the “Parareal”-based data showed significant performance improvements over E2E-JNet3 without extensive additional computational cost due to parallelization.

### 13 Feb 2024 v4

@@ 1 Introduction
Simulating wave propagation is an important task for many applications. For instance, when searching natural gas, wave propagation computations form the forward part of a numerical method for solving the inverse problem of geophysical inversion. By accurately analyzing the reflections and transmissions generated by complex media discontinuities, it becomes possible to characterize underground formations. However, traditional numerical computations often demand a computationally expensive fine grid representation to guarantee stability.
@@
$\mathfrak{u}_{n+1}\coloneqq\mathfrak{u}(x,t+\Delta t)=\mathcal{F}_{\Delta t}\mathfrak{u}_{n}\approx\Psi_{\Delta t}^{\theta}\mathfrak{u}_{n}\coloneqq\mathcal{I}^{\theta}\mathcal{G}_{\Delta t}\mathcal{R}\mathfrak{u}_{n}.$ (1)
@@ 2 Finite-Difference-Based Wave Propagators
As the two solvers operate on different Cartesian grids, with $\delta x<\Delta x$ and $\delta t<\Delta t$, we define the restriction operator $\mathcal{R}$, which transforms functions from a fine to a coarse grid, and the prolongation operator $\mathcal{I}$, which maps the inverse relation. The enhanced variants consist of (a) bilinear interpolations denoted as $\mathcal{R}$ and $\mathcal{I}^{0}$, while $\mathcal{I}^{0}\mathcal{R}\mathfrak{u}\neq\mathfrak{u}$, and (b) neural network components denoted as $\mathcal{I}^{\theta}\equiv\Lambda^{{\dagger}}\mathcal{I}_{\Delta t^{\star}}^{~{}\theta}\Lambda$, while the lower index indicates that the neural networks are trained when the step size $\Delta t^{\star}$ is used. For improved neural network inference, we use the transition operator $\Lambda$ to transform physical wave fields $(u,u_{t})$ to energy component representations $(\nabla u,c^{-2}u_{t})$, with $\Lambda^{{\dagger}}$ as the pseudo-inverse (see also nguyen2022numerical ; energy1 ). Figure 1 provides a schema visualizing the wave argument transitions.
Figure 1: Detailed schematic of E2E-JNet3 (adapted from nguyen2022numerical ). Each convolutional block (blue) encompasses a 3x3 convolutional layer ($\text{groups}=3$, $\text{padding}=1$), followed by a batch normalization and a ReLU activation function. To reduce the resolution, the 3rd and 6th convolutional block (red) employs a stride of two. Connectivity within the network is depicted by arrows, with the dashed arrow specifically indicating a single application of $\mathcal{G}_{\Delta t^{\star}}$.
@@
A simple model with bilinear interpolation (E2E-V, $\mathcal{I}^{0}\Psi_{\Delta t^{\star}}^{\theta}\mathcal{R}$) is used as a baseline. Each variant changes the baseline by exactly one aspect. This allows us to isolate the effect of each modification on the performance of the architecture.
The four investigated end-to-end models $\Psi_{\Delta t^{\star}}^{\theta}\coloneqq\mathcal{I}_{\Delta t^{\star}}^{~{}\theta}\mathcal{G}_{\Delta t^{\star}}\mathcal{R}$ are:
@@
The second baseline is taken from nguyen2022numerical and denoted as the modular, not end-to-end 3-level JNet (NE2E-JNet3), $\mathcal{I}_{\Delta t^{\star}}^{~{}\theta}(\mathcal{G}_{\Delta t^{\star}}\mathcal{R}$), while results of $\mathcal{G}_{\Delta t^{\star}}$ are used to separately train the E2E-JNet3 upsampling component.
@@ 3 Data Generation Approaches
    with $x\in[-1,1]^{2}$ and $\frac{1}{\sigma^{2}}\sim\mathcal{N}(250,10)$. $\tau\in[-0.5,0.5]^{2}$ represents the displacement of the Gaussian pulse’s location from the center.
@@
We sample $c\in\mathfrak{F}_{c}$ from randomly chosen subregions of two synthetic geological structures, Marmousi marmousi and BP bp , that are mapped onto the spatial grid $h\mathbb{Z}^{2}\cap[-1,1]^{2}$. For testing, four manually modified media are added to examine rapid variations in wave speed:
@@ 3.1 Multi-Step Training
$\mathfrak{u}_{n+k}\coloneqq\mathfrak{u}(t_{n}+k\Delta t)=(\mathcal{F}_{\Delta t})^{k}\mathfrak{u}_{n}\approx(\Psi_{\Delta t}^{\theta})^{k}\mathfrak{u}_{n}.$ (5)
@@ 3.2 Parareal Algorithm
Identical to nguyen2022numerical , our implemented scheme iteratively refines the solution using the difference between $\mathcal{F}_{\Delta t}$ and $\mathcal{G}_{\Delta t}$ for each subinterval $\Delta t$. In particular, missing high-frequency components occur due to the transition to a lower grid resolution, or a too simple numerical algorithm. Therefore, a more elaborate model $\Psi_{\Delta t}^{\theta}$ is required for convergence. Formally, we rearrange Eq. 1 for the time stepping of $\mathcal{F}_{\Delta t}$, and replace $\mathcal{F}_{\Delta t}(\mathcal{I}\mathcal{R})\mathfrak{u}_{n}$ by the computationally cheaper strategy $\Psi_{\Delta t}^{\theta}$ end-to-end:
@@
Parareal iterations alter a given initial sequence of wave fields $\mathfrak{u}_{n}^{0}$ to $\mathfrak{u}_{n}^{k}$ for $n\in\mathbb{N}$. This means that neural operators should be trained to map $\mathfrak{u}_{n}^{k}$ to $\mathcal{F}_{\Delta t}\mathfrak{u}^{k}_{n}$. Therefore, appropriate training patterns for this setup would naturally differ from those found in $\mathcal{D}$, and the dataset for use with Parareal should be sampled from a different distribution, denoted as $\mathcal{D}^{p}$.
@@ 4 Evaluation Setup
A. Comprehensive Training ($\mathcal{D}^{p}_{\text{train}}$): The models are trained according to the Parareal scheme in Eq. 7 and Eq. 8 with $K=4$ using a random sample that constitutes a quarter of the original dataset $\mathcal{D}$ for fair comparisons. The gradients are determined by summing the losses of a Parareal iteration.
@@ 5 Discussion
An overview of the upsampling architecture performances can be found in Table 1. As expected, the larger networks (E2E-Tira and E2E-Trans) performed slightly better compared to the 3-level JNet architecture, but for the ResNet architecture (E2E-JNet5), more weights did not increase accuracy by much. Consequently, we theorize that the ResNet design may be insufficient for capturing high-fidelity wave patterns, while especially highly-connected layers with an optimized feature and gradient flow (E2E-Tira) are better suited. Given that E2E-JNet3 ($\mathcal{D}^{w,m}$) had only a slightly worse average energy MSE on the test set, we generally advise against using the expensive models in our setup.
@@
While models trained with $\mathcal{D}^{p}_{\text{train}}$ have an unstable training progress and diverging loss, applying E2E-JNet3 ($\mathcal{D}^{p}_{\text{refine}}$) within the Parareal scheme showed better accuracy than E2E-JNet3 with Parareal (cf. Figure 5). As this training method improved the stability of Parareal, sampling the causality of concurrently solving multiple time intervals is an efficient enhancement to our end-to-end structure.
@@ 6 Conclusion
All end-to-end variants, including the variants with training modifications, outperformed the modular framework of nguyen2022numerical . In particular, the lightweight end-to-end 3-level JNet (E2E-JNet3) performed reasonably well given its low computation cost, and was further improved through a weighted, multi-step training scheme ($\mathcal{D}^{w,m}$) to feature time-dependent wave dynamics without adding complexity to the model or substantially extending the training duration. Similarly, the Parareal iterations using the neural propagator trained by the “Parareal”-based data showed significant performance improvements over E2E-JNet3 without extensive additional computational cost due to parallelization.